OpenStudy (anonymous):

The area of a rectangle is 450mm squared. The height is fifteen less than three times the base. What is the perimeter?

OpenStudy (anonymous):

we have the formula for the perimeter of a rectangle as 2(h+b) we know that h=3b-15 so perimeter = 2(h+b) = 2( (3b-15) + b) = 8b - 30 we are told that area = 450mm sqd area = b*h = b*(3b-15) = 3b^2 - 15b = 450mm using the quadratic formula we get 2 lengths for b, only one will make sense to use. Plugging in for the perimeter equation derived above should give you the answer